The owner of a car wash is trying to decide on the number of people to employ based on the following short-run production function:
Q = 6L – 0.5L2
Where Q = Number of car washers per hour
L = Number of workers
a. Generate a schedule showing total product, average product, and marginal product, Plot this schedule on a graph.
1
Expert's answer
2020-12-03T07:09:18-0500

a. The schedule can be calculated using next formulas: total product is "Q = 6L \u2013 0.5L^2" , average product is AP = Q/L = 6 - 0.5L, and marginal product is "MP = Q'(L) = 6 - L."